OpenStudy (anonymous):

solve: x^2 + 7x= -12 PLEASE HELP

jimthompson5910 (jim_thompson5910):

x^2 + 7x= -12 x^2 + 7x + 12 = 0 Find two numbers that multiply to 12 AND add to 7. They are: 4 and 3 So x^2 + 7x + 12 factors to (x+4)(x+3) which means x^2 + 7x + 12 = 0 turns into (x+4)(x+3) = 0

OpenStudy (anonymous):

wait you arent'suppose to supbtract 7x from that side and send it to the other side?

OpenStudy (anonymous):

OH WAIT NEVERMIND

jimthompson5910 (jim_thompson5910):

no, you want everything on one side

OpenStudy (anonymous):

mhmm and that's the final answer right?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

if (x+4)(x+3) = 0, then either x+4=0 or x+3=0 so x = ?? or x = ??

OpenStudy (anonymous):

the answer is 4 and 3 right?

jimthompson5910 (jim_thompson5910):

close, but not quite

OpenStudy (anonymous):

NOPE NEGATIVE

jimthompson5910 (jim_thompson5910):

yep -4 or -3

OpenStudy (anonymous):

OKAY THANKS :) while you're at it wanna answer another one :)

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

okayx^2 + 4x +4 = 36

jimthompson5910 (jim_thompson5910):

x^2 + 4x +4 = 36 x^2 + 4x +4 - 36 = 36 - 36 x^2 + 4x - 32 = 0 Find two numbers that multiply to -32 AND add to 4

OpenStudy (anonymous):

OKAY OKAY i get it thank you SOO much for the help! I really appreciate it.

OpenStudy (anonymous):

4 and pl

OpenStudy (anonymous):

hmm okay lemme try and figure this out. STAY ON THOUGH.

OpenStudy (anonymous):

me or the other guy?

OpenStudy (anonymous):

WHOW those are two different guys? lol you. since you started helping me with this

OpenStudy (anonymous):

i got it 8 and 4 right so 8-4 = 4x which is the middle correct?

OpenStudy (anonymous):

;D

OpenStudy (anonymous):

Ye[p

OpenStudy (anonymous):

No -8 4

OpenStudy (anonymous):

-8 & 4

OpenStudy (anonymous):

oh yeah since i need to get to -32

OpenStudy (anonymous):

don't u mean positive 8 and negative 4 LOL

OpenStudy (anonymous):

No x= -8 & x= +4

OpenStudy (anonymous):

Either: "complete the square" or use the "quadratic equation".

OpenStudy (anonymous):

Whichever you do... that is the answer. No lol.

OpenStudy (anonymous):

so it's x^2 +4x+4=36 x^2 +4x +4 -36+36-36 x^2 +4x-32=) (x+8)(x-4) x= -8,4 corret?

OpenStudy (anonymous):

Oui

OpenStudy (anonymous):

is that a yes?

OpenStudy (anonymous):

Ce

OpenStudy (anonymous):

i hope thats a yes. THANKS

OpenStudy (anonymous):

;D Yeess good luck